OpenStudy (anonymous):

Kitaer is a manager at a landscaping company. He has two workers to landscape an entire park, Cody and Kaitlyn. Cody can complete the project in 2 hours. Kaitlyn can complete the project in 1 hours. Kitaer wants to know how long it will take them to complete the project together.

OpenStudy (anonymous):

really confused on this?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x = time it takes for Cody to do the job alone y = time it takes for Kaitlyn to do the job alone z = time it take them to do the job together we're given x = 8 and y = 6. Solve for z \[\large \frac{1}{x}+\frac{1}{y}=\frac{1}{z}\] \[\large \frac{1}{8}+\frac{1}{6}=\frac{1}{z}\] \[\large \frac{3}{24}+\frac{4}{24}=\frac{1}{z}\] \[\large \frac{3+4}{24}=\frac{1}{z}\] I'll let you finish

OpenStudy (anonymous):

so 5/24?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

? Please explain

jimthompson5910 (jim_thompson5910):

\[\large \frac{3+4}{24}=\frac{1}{z}\] \[\large \frac{7}{24}=\frac{1}{z}\] cross multiply or take the reciprocal of both sides

OpenStudy (anonymous):

Oh! Okay :) Sorry, I didn't notice that

OpenStudy (anonymous):

so 7z=24

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

now fully isolate z

OpenStudy (anonymous):

so 24/7

jimthompson5910 (jim_thompson5910):

yep 24/7

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

Can I ask another please?

jimthompson5910 (jim_thompson5910):

sure, go ahead

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

what are some key features that you can think of? any at all?

OpenStudy (anonymous):

Maybe some degrees?

jimthompson5910 (jim_thompson5910):

you mean the degree of the polynomial?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what does the degree tell us?

OpenStudy (anonymous):

how many answers it would have

OpenStudy (anonymous):

I mean roots

jimthompson5910 (jim_thompson5910):

close, the degree tells us the maximum number of roots, or x-intercepts, possible example: x^3 + 7x^2 + 9 has at most 3 roots. It could have 3 roots, or 2 roots, or just 1 root.

OpenStudy (anonymous):

I understand. :)

jimthompson5910 (jim_thompson5910):

the degree also tells us how many turning points there are number of turning points = (degree) - 1 example: degree = 3 means we have 2 turning points